the amount of coffee cups sold daily at gloria jean's is uniformly distributed with a minimum of 100 cups and maximum of 600 cups. what is the probability that at least 300 cups will be sold.
Answer by mathmate(429)


Question:
the amount of coffee cups sold daily at gloria jean's is uniformly distributed with a minimum of 100 cups and maximum of 600 cups. what is the probability that at least 300 cups will be sold.

Solution:
Between the range of 100 and 600 cups, there are (600-100+1)=501 possible outcomes.
Since it is a uniform distribution, the probability of each possible outcome is 1/501.
Hence the probability of ≥ 300 cups is (600-300+1)/501=301/501.